OpenStudy (tabithax3):

I need help converting. A French automobile manufacturer claims that a sedan uses only 4.0 L of gasoline to travel 102.0 km. How many miles per gallon of gasoline could be expected from this car? (Note: 1 mile = 1.61 km)

OpenStudy (anonymous):

Convert 4 L to gallon. 1 liter = 0.264 gallon 4 liter = 4 X 0.264 = 1.056 gallon now convert 102.0 km into miles 1.61 km = 1 mile 102.0 km = 1/1.61 X 102.0 = 63.35 miles. Its mean 1.056 gallon is used to cover = 63.35 miles 1 gallon is used to cover = 63.35/1.056 X 1 = 59.99 or 60 miles per gallon.